# A Survey on Recent Advances in LLM-Based Multi-turn Dialogue Systems

### v1

@@ 6.2.2. Personalization and Consistency
Transitioning from the focus on dialogue personalization and consistency, Madotto et al. (madotto2019personalizing, ) use Model-Agnostic Meta-Learning (MAML) for personalizing dialogue learning. Their Persona-Agnostic Meta-Learning (PAML) framework treats different personas as separate tasks in meta-learning. PAML adapts dialogue models to new personas through dialogue samples, contrasting traditional persona-specific descriptions. This method is effective in generating fluent and consistent dialogues.
As the field advances towards more nuanced methods of personalization, Li et al. (li2021dialogue, ) have developed a Personalized Hybrid Matching Network (PHMN) that incorporates user-specific dialogue history into response selection. The PHMN model operates on two principal fronts: firstly, it extracts personalized wording behaviors from the user’s dialogue history. Secondly, it employs hybrid representation learning on context-response utterances, integrating a customized attention mechanism to extract essential information from context-response interactions. It improves the accuracy of matching responses to the user’s conversational style and preferences.
Building upon the concept of utilizing dialogue history for personalized response generation, Zhong et al. (zhong2022less, ) introduce the MSP model to refine user dialogue history. The MSP model includes User Refiner, Topic Refiner, Token Refiner, and Response Generator. The dialogue generation process integrates these components:
@@
where $\hat{y}$ is the response, $x$ the dialogue input, $u_{\text{sim}}$ the user similarity output, $t$ the topic information, and $A$ the token-level attention. This model enhances personalization in response generation by refining user dialogue history.
@@ 6.3. Hybrid Methods
Hybrid methods in ODD systems represent the integration of retrieval-based and generation-based approaches, combining the strengths of both. Early research in this area includes the work of Sordoni et al. (sordoni2015neural, ), who explored the use of context-sensitive generation models conditioned on traditional retrieval methods. This approach laid the groundwork for subsequent developments in the field. Another contribution by Yan et al. (yan2016docchat, ), introduced a model that dynamically selects between generating a response and retrieving one, depending on the conversation’s context. This blend allows for more flexible and contextually appropriate responses.
@@ 6.3.1. Integrating Retrieval and Generation
Building upon these initial explorations into hybrid methodologies, Borgeaud et al. (borgeaud2022improving, ) introduce the Retrieval-Enhanced Transformer (Retro) model, which combines a large-scale retrieval mechanism with an auto-regressive language model. This integration allows the language model to access contextually relevant information from a retrieval database. The model retrieval-enhanced sequence log-likelihood defined as:
@@
where $\mathbf{X}$ represents the input sequence, $\theta$ denotes the model parameters, $\mathcal{D}$ is the retrieval database, and $\text{Ret}\mathcal{D}$ refers to the retrieval operation.
@@ 6.3.2. Enhancing Dialogue with External Knowledge
Adolphs et al. ((adolphs2021reason, )) developed the Knowledge to Response (K2R) model, focusing on factual accuracy. K2R first generates knowledge sequences relevant to the dialogue context and then integrates this knowledge to synthesize the final response. Izacard and Grave ((izacard2020leveraging, )) developed the Fusion-in-Decoder (FiD) model, integrating generative models with passage retrieval. FiD encodes a question and multiple retrieved passages independently, concatenating each passage with the question for decoding, which facilitates the synthesis of information from various passages.
Expanding on the concept of retrieval augmentation, Xu et al. ((xu2021beyond, )) focus on the dynamics of long-term conversation. They explore Retrieval-Augmented Generative Models, characterized by their approach to integrating conversation context and model mechanics:
(33) $p(y|x)=\sum_{z\in\text{Retrieve}(x)}p(z|x)\cdot p(y|x,z).$
Additionally, they explored Memory-Based Models with Summarization, where the model summarizes past dialogues for response generation:
(34) $P(y|x,S)=\text{SummaryGeneration}(x,\text{Past Dialogues}).$
Singh et al. (singh2021end, ) develop $\mathrm{EMDR}^{2}$, integrating T5 encoding and decoding with a Top-k MIPS retrieval mechanism, enhanced by a scale evidence document encoder. The system aims to optimize multi-document reading and retrieval. $\mathrm{EMDR}^{2}$ implements an Expectation-Maximization (EM) algorithm for latent variable model training. The training process involves computing the posterior of the latent variable $Z$, and the training objective is given by:
(35) $L=\log p(a|q,Z_{\text{top-K}};\Theta)+\log\sum_{k=1}^{K}\text{SG}(p(a|q,z_{k};\Theta))p(z_{k}|q,Z_{\text{top-K}};\lambda),$
where SG is the stop-gradient operator, $\Theta$ are the reader’s parameters, and $\lambda$ are the retriever’s parameters.
Komeili et al. (komeili2021internet, ) introduce an ”Internet-Augmented Dialogue Generation” system, integrating real-time Internet. It consists of a search query generator and a response generation module. The search query generator, a transformer-based encoder-decoder model, generates internet search queries from the dialogue context. The response module uses the retrieved information to construct responses. The system’s function is encapsulated in:
(36) $R=\text{FiD}\left(\text{Encoder-Decoder}(C),\,\text{InternetSearch}(\text{Encoder-Decoder}(C))\right),$
where $C$ is the dialogue context and $R$ the generated response. The Encoder-Decoder function processes $C$, and InternetSearch retrieves information based on the query. The FiD method integrates this information to generate $R$.
@@ 7.1.1. Automatic Evaluation Approaches for Task-oriented Dialogue Systems
Joint goal accuracy (JGA), developed from Henderson et al. (henderson2014word, ) and Zhong et al. (zhong2018global, ) , is the most widely used evaluation approach for DST. The joint goal is a set of accumulated turn level goals up to a given turn in the dialogue. It compares the predicted dialogue states to the ground truth which includes slot values for all possible pairs. The output is considered as a correct prediction if and only if all the predicted values match its ground truth values at each turn. JGA can be expressed as:
@@
Slot accuracy (SA) (wu2019transferable, ) is also a widely used automatic evaluation approach. Unlike joint goal accuracy, it only compares each value to the corresponding ground truth individually without seeing other turns. SA can be expressed as:
@@
Average goal accuracy (AGA) (rastogi2020towards, ) is the average accuracy of predicting the correct value for an active slot in each turn. A slot becomes active if its value is mentioned in the current turn and is not inherited from previous turns. AGA can be expressed as:
@@
[…]
@@
BLEU (papineni2002bleu, ) is used to calculate the co-occurrence frequency of two sentences based on the weighted average of matched n-gram phrases. BLEU was originally used to evaluate machine translation, and has been used for evaluating TOD and ODD systems.
@@
Entity F1 is used to evaluated the model’s ability to generate relevant entities from the underlying knowledge base and to capture the semantics of the user-initiated dialogue flow. To compute an entity F1, one need to micro-average over the entire set of system dialogue responses and use the entities in their canonicalized forms.
@@ 7.1.2. Automatic Evaluation Approaches for Open-domain Dialogue Systems
Perplexity (vinyals2015neural, ) was originally conceived as an information-theoretic measure to assess how much a given language model is suited to predict a text sequence or, equivalently, how much a word sequence fits into a specific language model. It is now used as an analytic approach with potential application to support early diagnosis of symptoms of mental disorder. Perplexity can be expressed as:
@@
DIST-n (li2015diversity, ) is used to measure the diversity of response sequence for dialogue generation by calculating the number of distinct unigrams and bigrams in generated responses. The value is scaled by total number of generated tokens to avoid favoring long sentences. DIST-1 and DIST-2 are respectively the number of the distinct unigrams and bigrams divided by total number of generated words.
@@
Recall@K is one of the standard approaches to evaluate. For query $q$, it is defined as a radio of the number of relevant(positive) examples within the top-k ranked examples to the total number of relevant examples for $q$ given by $|P_{q}|$. It is denoted by $R^{k}_{\Omega}(q)$ when compute for query $q$ and database $\Omega$ and function $H(.)$ is the Heaviside step function. Therefore, it can be expressed as:
@@ 7.2. Human Evaluation
Human evaluation is also used as an evaluation approach in different tasks. Human evaluation focuses on the explanation of two matters: diversity and creativity, i.e., the capacity of varying their texts in form and emphasis to fit an enormous range of speaking situations, and the potential to express any object or relation as a natural language text. Furthermore, human evaluation scrutinizes three key aspects: Grammar (whether a generated sentence is grammatically correct and fluent), Faithful (whether the output accurately reflects the input), and Coherence (ensuring that a sentence is logically consistent and follows the natural flow of human writing).
@@ 8. DATASETS
In this section we introduce the datasets that have been widely used in TOD and ODD systems in recent years. Table 3. and Table 4. show some information for TOD and ODD datasets respectively.
@@ 8.1. Datasets for Task-oriented Dialogue Systems
Table 3. Overview of datasets for task-oriented dialogue. *SMD only provides statistics of avg of utterances per dialogue and avg of tokens per utterance. Single domain and multi domains show whether the dataset has dialogues that contain single domain or multi domains respectively.
@@
MultiWOZ 10,438 13.70 13.18 7
RiSAWOZ 11,200 13.57 10.91 12
CrossWOZ 6,012 16.90 16.25 5
P4G 1,017 10.43 - 1
WOZ 2.0 1,200 7.35 11.27 1
SMD 3,031 5.29* 9* 3
@@ 8.2. Datasets for Open-domain Dialogue Systems
Dailydialog (li2017dailydialog, ) is a multi-turn dialogue dataset with 13118 dialogues. It is created by scarping text from conversations held on an English learning website. The dialogues cover 10 topics and conform common dialog flows. Besides the dataset contains unique multi-turn dialog flow patterns, which reflect our realistic communication way. Each utterance is labeled with a dialogue act and an emotion.
@@
Table 4. Overview of datasets for open-domain dialogue. Human–Human denotes datasets where two people converse with each other. Scraped marks datasets which are gathered from an existing online resource.
@@
PersonaChat 164,356 Human-Human Crowdsourcing en
MMdialog 1,079,117 Scraped Social Media en
Dailydialog 13,118 Scraped - en
Pchatbot 198,875,796 Scraped Weibo, Judicial zh
PersonalDialogue $\approx$20.83 million Scaped Weibo zh
Douban 526,000 Scaped Douban zh
@@ 9. DISCUSSION
More and more researchers are investigating on applying LLMs to different components of a multi-turn TOD system or a muti-turn ODD system. An influential factor driving the popularity of multi-turn dialogue system tasks is the growing demand for chatbots in both industrial settings and everyday activities. Industry representatives such as OpenAI’s ChatGPT, Anthropic’s Claude 2, Meta’s Meta AI, and Google’s Gemini Ultra have greatly improved the convenience of people’s lives. Another reason is that a significant portion of natural language data exists in the form of dialogues, and multi-turn dialogue systems are more aligned with real-world scenarios, thus fostering the development of large model-based multi-turn dialogue systems. Meanwhile, LLMs provide humans with a powerful toolkit.
This section presents a discussion on challenges faced by LLM based multi-turn dialogue systems that deserved to be tackled and investigated future.
Deep understanding and Large-scale open-retrieval. By employing LLMs in multi-turn conversations, it requires a more proficient understanding and retention of longer contextual information to generate responses that are more coherent and relevant. In contrast to open-domain QA tasks, open-retrieval dialogues pose new challenges primarily arising from the features of human-machine interaction, impacting both efficiency and effectiveness.
Emotionalization and Personalization. Emotionalizing dialogue systems and enhancing logical consistency can make responses more aligned with the needs of the queries, understand the emotions behind expressions, thus better understanding the semantic and contextual connections, and providing more appropriate answers.
Multi-task dialogue systems. Recent research in end-to-end TOD systems and knowledge-grounded open-domain systems has opened up the prospect of merging these two distinct paradigms within a cohesive framework, or perhaps even a unified model. These hybrid dialogue systems are designed to operate concurrently as both assistants, efficiently executing specific tasks, and as chatbots, engaging in conversational interactions.
Multi-model dialogue systems. The world we inhabit is inherently multi-modal, with humans utilizing a multitude of senses such as sight, hearing, smell, taste, and touch to perceive it. Hence, it is imperative that chatbots possess the proficiency to blend information from disparate modalities. While contemporary large chat models exhibit admirable effectiveness in processing text, audio, and pictures, they still encounter significant limitations and challenges when it comes to video processing.
Bias identification and Privacy protection. LLMs have the potential to generate content that is harmful, offensive, or biased due to their training on publicly available online datasets. While researchers have endeavored to address this through fine-tuning, challenges may persist, particularly in languages other than English where publicly available datasets are limited. It is important to prioritize user privacy alongside the advancement of more sophisticated dialogue systems.

### v2

@@ 4.2.2. Personalization and Consistency
Madotto et al. (madotto2019personalizing, ) proposed the Persona-Agnostic Meta-Learning (PAML) framework, applying Model-Agnostic Meta-Learning to personalized dialogue generation. This approach treats individual personas as separate learning tasks, enabling the model to adapt to new personas through limited dialogue samples rather than requiring explicit persona descriptions. PAML implements a meta-learning objective that optimizes for quick adaptation to new personas while maintaining general dialogue generation capabilities.
Li et al. (li2021dialogue, ) developed the Personalized Hybrid Matching Network (PHMN), focusing on integrating user-specific dialogue history into response selection. PHMN operates through dual pathways: first extracting personalized linguistic patterns from historical dialogues, then implementing a hybrid representation learning mechanism with customized attention for context-response matching. The model processes utterance-response pairs through multiple interaction layers with self-attention mechanisms, enabling it to capture both local and global personalization patterns.
Zhong et al. (zhong2022less, ) introduced the MSP model, implementing a multi-component architecture for dialogue history refinement. The system consists of specialized modules including User Refiner, Topic Refiner, Token Refiner, and Response Generator, each serving specific functions in the personalization process. The dialogue generation process integrates these components according to:
@@
where $\hat{y}$ represents the generated response, $x$ denotes the dialogue input, $u_{\text{sim}}$ indicates user similarity metrics, $t$ provides topic-related information, and $A$ represents token-level attention patterns. The model implements iterative refinement of user dialogue history to enhance response personalization and maintains consistency through continuous updates to user and topic representations.
These generation-based methods collectively address fundamental challenges in ODD systems through various technical innovations. Knowledge enhancement techniques enable models to access and incorporate external information for more informed and accurate responses. Persona management mechanisms implement sophisticated methods for maintaining consistent user interactions across extended conversations. Recent advances in self-reflection and critique mechanisms, particularly demonstrated by Self-RAG, provide frameworks for more controlled, verifiable, and adaptable response generation. The integration of these approaches continues to advance the capabilities of dialogue systems while presenting new opportunities for improved human-machine interaction.
@@ 4.3. Hybrid Methods
Research in open-domain dialogue systems has explored methods combining retrieval-based and generation-based approaches to address their respective limitations. Retrieval-based methods can provide factual and grounded responses but are constrained by their predefined response sets. Generation-based methods offer flexibility but may produce inconsistent or unfactual responses. Early work by Sordoni et al. (sordoni2015neural, ) established context-sensitive generation models conditioned on retrieved results, demonstrating the feasibility of combining both approaches. Yan et al. (yan2016docchat, ) developed a system that dynamically alternates between retrieval and generation based on dialogue context requirements, introducing adaptive mechanisms for response selection.
@@ 4.3.1. Integrating Retrieval and Generation
Recent research has focused on tighter integration between retrieval and generation components. Borgeaud et al. (borgeaud2022improving, ) introduced the Retrieval-Enhanced Transformer (Retro) model, which incorporates large-scale retrieval mechanisms within the transformer architecture. The model implements retrieval-enhanced sequence log-likelihood according to:
@@
where $\mathbf{X}$ represents the input sequence, $\theta$ denotes model parameters, $\mathcal{D}$ indicates the retrieval database, and $\text{Ret}\mathcal{D}$ refers to the retrieval operation. This formulation enables the model to condition its generation on retrieved contextual information at multiple sequence positions.
Yu et al. (yu2023chain, ) developed Chain-of-Note (CON), addressing key challenges in retrieval-augmented language models through sequential reading notes. The framework implements three primary components: a retrieval mechanism determining when additional knowledge is required, an evaluation module assessing retrieved information relevance, and a generation module producing responses with self-critiques. CON introduces specialized mechanisms for handling different types of retrieved information, including processes for relevant document integration, contextual inference from partially relevant information, and unknown response detection. The framework also implements adaptive retrieval thresholds and noise handling procedures to maintain generation quality under varying retrieval conditions.
@@ 4.3.2. Enhancing Dialogue with External Knowledge
The integration of external knowledge sources has emerged as a key direction in hybrid dialogue systems. Adolphs et al. (adolphs2021reason, ) proposed the Knowledge to Response (K2R) model, which first generates knowledge sequences relevant to the dialogue context and then integrates this knowledge during response generation. This two-stage process enables the model to maintain factual accuracy while producing contextually appropriate responses. Izacard and Grave (izacard2020leveraging, ) developed the Fusion-in-Decoder (FiD) model, implementing independent encoding of questions and multiple retrieved passages before concatenated decoding, facilitating efficient information synthesis from diverse sources.
@@ 5.1.1. Automatic Evaluation Approaches for Task-oriented Dialogue Systems
Joint goal accuracy (JGA), developed from Henderson et al. (henderson2014word, ) and Zhong et al. (zhong2018global, ) , is the most widely used evaluation approach for DST. The joint goal is a set of accumulated turn level goals up to a given turn in the dialogue. It compares the predicted dialogue states to the ground truth which includes slot values for all possible pairs. The output is considered as a correct prediction if and only if all the predicted values match its ground truth values at each turn. JGA offers a stringent evaluation of the model’s capacity to capture all dialogue states, ensuring high reliability. However, it does not accommodate natural dialogue variations and may overlook subtle shifts in user intentions. JGA can be expressed as:
@@
Slot accuracy (SA) (wu2019transferable, ) is also a widely used automatic evaluation approach. Unlike joint goal accuracy, it only compares each value to the corresponding ground truth individually without seeing other turns. SA is a straightforward metric for evaluating individual slot predictions, but it disregards contextual information from previous turns and long-range dependencies, limiting its effectiveness in complex dialogues. SA can be expressed as:
@@
Average goal accuracy (AGA) (rastogi2020towards, ) is the average accuracy of predicting the correct value for an active slot in each turn. A slot becomes active if its value is mentioned in the current turn and is not inherited from previous turns. AGA emphasizes predicting active slots, providing a practical gauge of system responsiveness. However, it overlooks the broader dialogue context, essential for tasks requiring long-term understanding. AGA can be expressed as:
@@
Requested slot F1 indicates the model performance in correctly predicting if a requested slot is requested by the user, estimated as the macro-averaged F1 score over for all requested slot. The macro-averaged of F1 score is computed over the individual slot-type and slot-value for every turn. Requested Slot F1 evaluates the model’s ability to predict requested slots accurately, crucial for interpreting user requests. However, it may exhibit bias towards certain slot types and struggle with balancing precision and recall. To define the macro-averaged F1 score($ma\,F_{1}$), first consider the following precision($P_{i}$) and recall ($R_{i}$) within each class, i= 1,, 2i\,=\,1,\ $\dots$,\ 2:
@@
BLEU (papineni2002bleu, ) is used to calculate the co-occurrence frequency of two sentences based on the weighted average of matched n-gram phrases. BLEU was originally used to evaluate machine translation, and has been used for evaluating TOD and ODD systems. While effective for measuring fluency via n-gram matching, it overlooks semantic meaning and contextual coherence, limiting its ability to assess dialogue depth and naturalness.
@@
Entity F1 is used to evaluated the model’s ability to generate relevant entities from underlying knowledge base and to capture the semantics of the user-initiated dialogue flow. To compute an entity F1, one need to micro-average over the entire set of system dialogue responses and use the entities in their canonicalized forms. Entity F1 is crucial for entity recognition but emphasizes precision over fluency, neglecting dialogue consistency and coherence.
@@ 5.1.2. Automatic Evaluation Approaches for Open-domain Dialogue Systems
Perplexity (vinyals2015neural, ) was originally conceived as an information-theoretic measure to assess how much a given language model is suited to predict a text sequence or, equivalently, how much a word sequence fits into a specific language model. It is now used as an analytic approach with potential application to support early diagnosis of symptoms of mental disorder. Perplexity quantifies the model’s ability to predict a word sequence, but it fails to capture semantic coherence or the overall quality of the dialogue, limiting its applicability for evaluating natural, contextually-rich conversations. Perplexity can be expressed as:
@@
DIST-n (li2015diversity, ) is used to measure the diversity of response sequence for dialogue generation by calculating the number of distinct unigrams and bigrams in generated responses. The value is scaled by total number of generated tokens to avoid favoring long sentences. DIST-1 and DIST-2 are respectively the number of the distinct unigrams and bigrams divided by total number of generated words. While DIST-n encourages diversity, it may compromise coherence and relevance in favor of generating more varied responses.
@@
Recall@K is one of the standard approaches to evaluate. While it is effective for evaluating relevance, it does not provide an assessment of the quality of individual responses. For query $q$, it is defined as a radio of the number of relevant(positive) examples within the top-k ranked examples to the total number of relevant examples for $q$ given by $|P_{q}|$. It is denoted by $R^{k}_{\Omega}(q)$ when compute for query $q$ and database $\Omega$ and function $H(.)$ is the Heaviside step function. Therefore, it can be expressed as:
@@ 5.2. Human Evaluation
Human evaluation is also used as an evaluation approach in different tasks. Human evaluation focuses on the explanation of two matters: diversity and creativity, i.e., the capacity of varying their texts in form and emphasis to fit an enormous range of speaking situations, and the potential to express any object or relation as a natural language text. Furthermore, human evaluation scrutinizes four key aspects: Grammar (whether a generated sentence is grammatically correct and fluent), Faithful (whether the output accurately reflects the input), Coherence (ensuring that a sentence is logically consistent and follows the natural flow of human writing), and Consistency(ensuring that the dialogue remains stable over multiple turns).While these metrics are crucial, excessive emphasis on any one can introduce limitations. Over-focusing on grammar and faithfulness may constrain creativity and diversity in responses, while an overemphasis on coherence can reduce flexibility, resulting in rigid or overly predictable dialogues. Similarly, prioritizing consistency may hinder the model’s adaptability to context shifts or new topics.
@@ 6. DATASETS
In this section we introduce the datasets that have been widely used in TOD and ODD systems in recent years. Table 4. and Table 5. show some information for TOD and ODD datasets respectively. Examples of four commonly used dialogue datasets are shown in Figure 5.
Figure 5. Examples of four widely-used datasets. Some text is shortened for space.
@@ 6.1. Datasets for Task-oriented Dialogue Systems
Table 4. Overview of datasets for task-oriented dialogue. *SMD only provides statistics of avg of utterances per dialogue and avg of tokens per utterance. Single domain and multi domains show whether the dataset has dialogues that contain single domain or multi domains respectively. You can download the datasets by clicking the blue text.
@@
MultiWOZ(budzianowski2018multiwoz, ) 10,438 13.70 13.18 7
RiSAWOZ(quan2020risawoz, ) 11,200 13.57 10.91 12
CrossWOZ(zhu2020crosswoz, ) 6,012 16.90 16.25 5
P4G(wang2019persuasion, ) 1,017 10.43 - 1
WOZ 2.0(mrkvsic2016neural, ) 1,200 7.35 11.27 1
SMD(eric2017key, ) 3,031 5.29* 9* 3
@@ 6.2. Datasets for Open-domain Dialogue Systems
Dailydialog (li2017dailydialog, ) is a multi-turn dialogue dataset with 13118 dialogues. The dialogues cover 10 topics and conform common dialog flows. Besides the dataset contains unique multi-turn dialog flow patterns, which reflect our realistic communication way. Each utterance is labeled with a dialogue act and an emotion.
@@
Table 5. Overview of datasets for open-domain dialogue. Human–Human denotes datasets where two people converse with each other. Scraped marks datasets which are gathered from an existing online resource. You can download the datasets by clicking the blue text. Here, en denotes the English dataset and zh denotes the Chinese dataset.
@@
PersonaChat(zhang2018personalizing, ) 164,356 Human-Human Crowdsourcing en
MMdialog(feng2022mmdialog, ) 1,079,117 Scraped Social Media en
Dailydialog(li2017dailydialog, ) 13,118 Scraped - en
Pchatbot(qian2021pchatbot, ) 198,875,796 Scraped Weibo, Judicial zh
PersonalDialogue(zheng2019personalized, ) $\approx$20.83 million Scaped Weibo zh
Douban(wu2016sequential, ) 526,000 Scaped Douban zh
@@ 7. DISCUSSION
The rapid advancement of LLMs has revolutionized TOD and ODD systems, particularly in multi-turn interactions. While numerous studies have focused on leveraging LLMs to improve response coherence, fluency, and relevance, challenges persist that limit their applicability in real-world, sustained dialogue scenarios. This section introduce the challenges and future directions for LLM-based multi-turn dialogue systems.
High Computational Resource Requirements. Multi-turn conversations typically involve long contextual dependencies, introducing significant reasoning overhead. LLMs demand substantial computing power for both training and inference. For instance, training models such as Megatron-LM requires thousands of A100 GPUs, rendering it impractical for use in resource-constrained environments (samsi2023words, ). Although model compression techniques such as quantization and knowledge distillation (lin2024awq, ; tian2025knowledge, ) have been proposed to mitigate this issue, they frequently lead to performance degradation. Without significant advancements in either hardware efficiency or algorithmic innovation, the computational cost will remain a critical bottleneck to the scalable deployment of LLM-based dialogue systems.
Looking forward, we believe that the key breakthrough will come from designing dialogue-specific architectures that balance contextual reasoning with efficiency. Instead of one-size-fits-all LLMs, future systems may adopt hybrid strategies—leveraging smaller, distilled models for routine dialogue turns while invoking powerful LLMs for complex reasoning when necessary. Additionally, continual learning on edge devices, supported by efficient parameter updates (e.g., LoRA, adapters), may offer a viable pathway for low-resource deployment.
Long-Context Modeling. Long-context modeling remains a major challenge in multi-turn dialogue systems. Due to the fact that multi-turn tasks typically involve extensive dialogue history, the context will inevitably contains large amounts of lengthy information. Despite many approaches attempt to address this issue by compressing dialogue history or by extending the model’s context window to preserve contextual coherence across longer interactions (shin2022dialogue, ; feng2024fact, ; shang2025longrope2, ), LLMs still struggle to maintain consistent discourse across extended interactions. They may forget, overlook, or misrepresent previous dialogue turns, resulting in incoherent, contradictory, or repetitive responses (li2025beyond, ). This exposes fundamental limitations of the Transformer architecture in modeling long-range dependencies.
From our perspective, long-context modeling in dialogues is not merely a memory challenge, it is a representation learning problem. Effective dialogue systems must be able to dynamically abstract and selectively attend to salient past turns based on intent and role-specific context. Techniques like structured memory (e.g., memory graphs), dialogue act tracking, and adaptive compression schemes may provide more cognitively plausible solutions than raw token window extension. Moreover, incorporating human-like ”summary turns” or retrospection prompts within interactions could help models preserve narrative coherence across long conversations.
Domain Adaptation and Robustness. LLMs often exhibit degraded performance when handling specialized domains (e.g., medical, legal) or noisy inputs. However, multi-turn dialogue systems, especially TOD systems, typically need to operate across diverse domains, making the cost of building fully LLM-based multi-turn dialogue systems prohibitively high.
Current research strategies primarily involve either training LLMs or employing few-shot ICL to complete multi-domain dialogue tasks (yang2025omnidialog, ; yi2024intent, ). Training offers a potential remedy by adapting LLMs to target domains, but it can compromise model robustness, generalization and security. Training methods often exhibit poor practical usability due to its limited training data. They will also reduce resistance of LLMs to adversarial attacks. Furthermore, even slight perturbations in input can produce significantly different outputs, highlighting fragility in generalization. Few-shot ICL, meanwhile, yields unstable performance and introduces additional conversational context, complicating dialogue history. Consequently, future work must explore principled approaches that enable secure and robust domain adaptation without degrading overall performance or safety.
Multi-turn Interactions of Agentic AI. Multi-turn dialogue is also an important research domain for LLM agents, as it is an essential aspect of their capabilities. Moreover, the combination of agent’s external tool calling and multi-turn dialogue process will enhance the rationality and richness of responses. In particular, TOD systems routinely call external APIs to assist users with flight search and booking: querying real‑time schedules, checking seat availability, comparing fares across airlines, and completing reservations. By grounding each turn in up-to-date information from external tool calls, multi-turn dialogue systems can provide more accurate and comprehensive knowledge, better meeting user needs and accommodating personalized preferences during conversations.
Textual agents such as ChatDev (chatdev, ) and MetaGPT (metagpt, ) construct heterogeneous teams of role‑specialized agents to tackle complex tasks, but these approaches still exhibit limitations in their ability to manage external tool invocations. We posit that enabling LLM‑driven agents to interact directly with external environments unlocks new potential for handling multi-turn dialogue tasks (agentbench, ). Recent agent frameworks demonstrate tool‑use capabilities, ranging from invoking code compilers to interacting with web interfaces, which can be grafted onto multi‑turn dialogue models (lemur, ; alfworld, ; mind2web, ). By equipping TOD systems with these advanced agent techniques for external tool calling, we can substantially enhance their ability to engage with the real world and fulfill ever more complex user requests.
Privacy Protection. Privacy protection in dialogue systems has become an increasingly urgent issue as these systems frequently process highly sensitive user data. The increasing prevalence of adversarial prompt-injection and prompt-leaking attacks has demonstrated that LLM-driven multi-turn dialogue systems can unintentionally expose private information or internal prompts, posing serious privacy risks (agarwal2024prompt, ). Moreover, unlike static text, dialogues present unique privacy challenges that users will reveal personal information continuously and implicitly. This reality underscores the pressing need for robust privacy-preserving mechanisms within interactive dialogue settings, especially as these systems proliferate in consumer-facing and enterprise environments.